OpenStudy (anonymous):

You completed 12 laps on a running track in 25 minutes to the nearest minute how long will it take you to complete a total of 20 laps

OpenStudy (anonymous):

what do you think the ratio will be?

OpenStudy (anonymous):

12/25=x/20

OpenStudy (anonymous):

is this correct

OpenStudy (anonymous):

no, you flipped the second fraction again. since you have laps/minutes first, your second fraction should be laps/minutes

OpenStudy (anonymous):

oh so 12/25 = 20/x

OpenStudy (anonymous):

ok its right thanks

OpenStudy (anonymous):

so what would it be

OpenStudy (anonymous):

41.6

OpenStudy (anonymous):

as long as that's 500/12

OpenStudy (anonymous):

ya so 41.6 minuttes

OpenStudy (anonymous):

yep